OpenStudy (anonymous):

Find 3 consecutive even integers such that the square of the largest exceeds the sum of the quarters of the squares of the other two by 12

OpenStudy (mathteacher1729):

Wow, this is a complex statement. How do you feel with simpler problems of this nature? "The sum of three consecutive integers is 12" for example. or "The sum of two consecutive odd numbers is 16" There is quite a bit to break apart in this prob, that's why I ask.

OpenStudy (anonymous):

I am lost on integers

OpenStudy (mathteacher1729):

Oh, ok. Then any explanation of this problem will probably seem way too confusing. Let's start with the ones I wrote instead. Those are a bit simpler, and will help build up to this problem.

OpenStudy (mathteacher1729):

An integer is a kind of number. It's a whole number which can be positive or negative. Also, zero is an integer. ... -3 , -2, -1, 0, 1, 2, 3, ... are all integers.

OpenStudy (mathteacher1729):

Do you know what "consecutive" means?

OpenStudy (anonymous):

yes in a row ie 12 13 14

myininaya (myininaya):

x is the first even integer x+2 is the next even integer x+4 is the third even integer (x+4)^2=1/4*(x^2+(x+2)^2)+12

myininaya (myininaya):

x^2+8x+16=1/4*(x^2+x^2+4x+4)+12 x^2+8x+16=1/4*(2x^2+4x+4)+12 x^2+8x+16=x^2/2+x+1+12 x^2+8x+16=x^2/2+x+13 -x^2/2+7x+3=0 x^2-14x-6=0 this does not give us integer so maybe i misread the question

OpenStudy (anonymous):

To me it seems many of these questions are not well written which doesn't help

myininaya (myininaya):

the question seems really confusing lol

OpenStudy (anonymous):

lol tell me about it. They are ALL that way

OpenStudy (mathteacher1729):

Myininaya -- what if we start with x = 1 :-p You need to do 2n , 2n+2 , 2n+4. Also, this is a complex problem. Is this a test or HW question?

OpenStudy (mathteacher1729):

It drives me mad when profs put questions on the test which are WAY MORE DIFFICULT and NOTHING LIKE the HW problems they assign or the problems they go over in class.

OpenStudy (anonymous):

yeah, thanks for help. Is homework

myininaya (myininaya):

it doesn't matter what kindof question this is this is an algebra class lol this is ridiculous for an algebra class

OpenStudy (anonymous):

THANK YOU!!! I agree

myininaya (myininaya):

unless it is a bonus question

OpenStudy (mathteacher1729):

Well in the proper context, it's not a ridiculous question. If this was a question at the end of the chapter on how to interpret word problems for an advanced algebra class, yeah it makes sense. Is this a TYPICAL hw problem? Like, are all the others similar in complexity to it?

OpenStudy (anonymous):

This is PRE algebra. They are all this way

OpenStudy (anonymous):

and not a bonus question regular one

myininaya (myininaya):

i haven't gave up yet you wrote the question exactly as it appears in the book right?

OpenStudy (anonymous):

yes

myininaya (myininaya):

others or other?

myininaya (myininaya):

nvm

OpenStudy (anonymous):

I cant wait for this class to be over

myininaya (myininaya):

i don't see how it isn't (x+4)^2=1/4*x^2+1/4*(x+2)^2+12 x^2+8x+16=1/4*x^2+1/4*(x^2+4x+4)+12 x^2+8x+16=1/4*x^2+1/4*x^2+x+1+12 x^2+8x+16=1/2*x^2+x+13 1/2*x^2+7x+3=0 x^2+14x+6=0 but like i said before this does not give an integer for x so i just don't get the question i guess mathteacher will you please tell me what is wrong with this

OpenStudy (anonymous):

I appreciate your help

myininaya (myininaya):

i tried lol i can't see this problem any other way what i have above is what it means to me maybe i'm making a mistake in the translation, but i don't see where or how so sorry

OpenStudy (anonymous):

No need to be sorry, I appreciate it.

OpenStudy (mathteacher1729):

Ok, got something, but I think there might be an issue with the wording here. Lemmie know what you think.

myininaya (myininaya):

that is exactly how i set my up exept we introduced are variables differently

OpenStudy (anonymous):

makes more sense than what I got

myininaya (myininaya):

hey that is 12 right not 15?

OpenStudy (anonymous):

yes

myininaya (myininaya):

ok i have discussed this problem with others and they also agree there are no such integers i'm just shocked they would have given you a problem that you cannot find the solution to so just say there are no such integers

OpenStudy (anonymous):

ok thank you